The total cost for making umbrellas for Ella, Inc. is given by the equation TC= q2+1/(3+q)+3. What is their fixed cost?

a.4

b.3.33

c.4.25

d.3
Answer #1

Answer

fixed cost is the total cost when q=0

TC=q^2+1/(3+q)+3

=0^2+1/(3+0)+3

=1/3+3

=0.33+3

=3.33

the fixed costs are $3.33

option b